OpenStudy (miss.rose):

Question 4 of Triangle help!

rebeccaxhawaii (rebeccaxhawaii):

mhm

rebeccaxhawaii (rebeccaxhawaii):

36+56+?=180

OpenStudy (jdoe0001):

hint: sum of all internal angles in a triangle is 180

OpenStudy (miss.rose):

36+56= 92 180-92= 88?